# Tenspiler: A Verified Lifting-Based Compiler for Tensor Operations

### v1

@@ 5 Synthesis Optimizations
Figure 6: A depth 4 general synthesis grammar for the source in Figure 1(a)
@@ 5.1 Restricting Operators
all reduction operations are excluded. In Fig. 6, the operators in $v_{4}$ and $l_{4}$ will be removed (shown in red). These correspond to operators returning 1-D vectors and integers respectively.
@@ 5.2 Restricting Program States
We further optimize the search space by restricting the set of program states in Eq. 2. Instead of satisfying the VC for all $\sigma$, we find $ps$ and $inv$ that satisfy a bounded set. Bounded synthesis is crucial because most SyGuS solvers have limited support for recursive function definitions and require SMT solvers for validation. However, SMT solvers lack inherent support for reasoning about Tensir operators that are not covered by the standard theories defined in SMT-LIB [9] and require additional axioms to be defined. We instead integrate bounded synthesis by restricting the maximum unrollings of recursive operators, thereby eliminating the need for additional axioms. Specifically, we restrict the program states by limiting the lengths of the data structures and the sizes of the data types. For instance, in Tenspiler, we constrain all 1D tensors to length 2 and the integers to 6 bits or less for the first rounds of synthesis. In Fig. 6, all the tensor literals in $t_{1}$ are changed from an unbounded length “n” (shown in orange) to length 2 (shown in blue). If the synthesized choices fail to verify, we then increase the bounds in subsequent rounds. Note that since the synthesized solutions only work for a restricted set of program states, we invoke the theorem prover for subsequent verification to check if $ps$ and $inv$ are valid for all states.
@@ 5.3 Leveraging Expression Trees
Despite the above two optimizations, the synthesis search space remains large. For example, in the context of S in Figure 1(a) for which we need to synthesize two $inv$s and one $ps$, a depth-4 grammar, after removing the reduction operations, still presents around 100k potential solutions just for $ps$. Tensir plays a significant role in the further pruning of this search space. The design of Tensir operators effectively bridges the gap between high-level tensor operations and the loop-based paradigm commonly used for computing on tensors. This property of Tensir allows us to leverage the expression tree-based filtering approach, which we describe next, to efficiently prune the synthesis search space.
@@ 5.4 Constraining Variables
The final optimization is to
pinpoint specific variables (vars such as a,b in Figure 1(a)) and literals (lits such as 255 in Figure 1(a)) to be used in the grammar. Specifically, we constrain the variables to the set of live variables and also constrain constants to the set of constants that have appeared in the program. This strategy simplifies the computational task, avoiding the complexity of synthesizing a complete depth-4 operator sequence. By leveraging our expression tree-based approach, the synthesizer promptly yields the correct solution within 76 secs: t_t(t_t(b, a, +), t_s(t_t(b, a, *), 255, /), -).
@@ 5.5 Overall Synthesis Algorithm
The algorithm described in Fig. 7 summarizes the synthesis phase in Tenspiler. This phase is used to search for the bodies of PS and invariants which satisfy the VC. The synthesis is an iterative process conducted over multiple rounds, assuming a filtered search space leveraging type-based and expression tree
optimizations described earlier. We start with the tensor bound size set to 2 (line 2) which corresponds to the restricting the program states optimization. In each round, we invoke Rosette’s search algorithm (line 5) to generate candidates for $ps$ and $invs$. Upon obtaining a solution, the candidate undergoes validation against the VC for all program states, as the synthesis phase only checks within a constrained set of program states. We invoke a verifier (CVC5) (line 9) to perform this check. If the verifier yields “UNSAT”, the generated candidates are correct. Conversely, if “SAT” is returned, indicating incorrect candidates (line 11), the VC is augmented with blocking constraints. These constraints state that the generated $ps$ or $inv$ in next round should differ from those in the previous rounds. This iterative process continues for a specified number of rounds (max_rounds) before incrementing the tensor bound sizes. In cases where Rosette’s search algorithm does not produce a solution initially (line 13), indicating an overly restrictive grammar, the initial grammar is expanded to include additional options for both PS and invariants, such as choices for loop bounds, indexing, and operator sequences. We keep a separate timer (not shown in Fig. 7) that maintains a maximum time bound for the entire synthesis process.
1def synthesis_algorithm(spec, holing_grammar, search_algorithm, verifier, max_rounds, timeout):
2 r, k = 0, 2 # rounds within one list bound, tensor size bound
@@
8 if verifier(specification, ps, inv) == "UNSAT": return ps, inv
@@
14 k += 1
15 return None # No solution found
@@ 6 Experiments
    simpl_array: 5 functions for computations on integer arrays originally from the work of So and Oh [33].
@@ 6.1.1 Datasets for Evaluation
For the Llama benchmarks, we evaluate the synthesized code using weights from Vicuna [16], an open-source LLM with similar model size as Llama2666We did not use Llama2 model weights as they are not publicly available.. Some kernels operate over model weights, for which we directly use the weight matrices from Vicuna. For kernels operating over inputs, we simulate embeddings by creating random 32-bit float vectors within the range [0, 1). The evaluation primarily uses the 33B-parameter Vicuna model; however, for evaluating the MLX framework, the 7B-parameter version is used due to memory limitations.
@@ 6.1.2 Target Software Frameworks and Hardware Accelerators
To establish a baseline for execution performance, we compile C++ code for all the benchmarks using gcc-8.3.0 with -O3 flag and then run them on an Intel Xeon 8380 CPU. Given that each DSL is tailored to enhance performance on specific hardware, we conduct evaluation across five distinct platforms: the Intel Xeon 8380 CPU, Nvidia V100 GPU, Apple M1 Pro, Intel Gaudi2 processor, and the Gemmini accelerator [^7] Due to lack of physical device, Gemmini evaluations are done on a simulator with limited computing power and no file system support. Thus, it is compared with smaller random inputs.. In all, we utilized 7 different DSL-hardware device combinations for our experiments: (1) NumPy-CPU, (2) TensorFlow-V100, (3) PyTorch-V100, (4) MLX-Apple M1 (5) TPCC-Gaudi, (6) PyTorch-Gaudi, and (7) Gemmini. This comprehensive mapping of each backend to its corresponding device enables us to accurately assess the benefits Tenspiler provided through lifting.
The exact configurations of all the hardware devices and their software environments are described in the (Appendix C) due to lack of space.
@@ 6.2 Synthesis Timings
In this section, we present the time Tenspiler takes to synthesize and verify solutions for each of our benchmarks. During the synthesis phase, we apply all the optimizations mentioned in Sec. 5 with a 1-hour timeout. Tenspiler synthesizes the correct translations for all the benchmarks under 15 mins.
(a) Synthesis timings for the blend benchmarks (image processing)
(b) Synthesis timings for the utdsp benchmarks
(c) Synthesis timings for the DSP benchmarks
(d) Synthesis timings for the simpl_array benchmarks
(e) Synthesis timings for the darknet benchmarks
(f) Synthesis timings for the blas benchmarks
(g) Synthesis timings for the mathfu benchmarks
(h) Synthesis timings for the dspstone benchmarks
(i) Synthesis timings for the Llama benchmarks (ML kernels)
(j) Synthesis timings for the makespeare benchmarks
Figure 8: Synthesis timings for all the benchmark suites. Benchmark name legend in  Appendix B.
@@ 6.2.1 Analysis
We recognize that the tree approach may restrict the ability of Tenspiler to synthesize different solutions. However, through manual verification, we confirm that Tenspiler consistently generates optimal solutions across our benchmarks. We evaluate the solutions using expression length as the cost function. Generally, shorter expressions mean fewer function calls and thus lower execution cost. To illustrate, we use the linear_dodge example from the blend benchmarks for which the synthesized solution is as follows:
@@
Figure 9: Kernel speedup over baseline. Benchmark legend in  Appendix B.
@@ 6.3 Performance Timings
Figure 10: E2E speedup over baseline. Benchmark legend in  Appendix B.
@@ 6.4 Ablation Study
Figure 11: Synthesis timings for all the benchmarks with and without Tenspiler’s tree-based optimization. Benchmark name legend in  Appendix B.
@@
As illustrated in Fig. 11, without the tree based optimization, 42 out of the total 69 benchmarks timed out. In particular, all benchmarks of the blend, blas, dspstone, makespeare, and utdsp suites timed out. For the benchmarks that succeed, the synthesis phase slows down by an average 101.55$\times$ compared to the tree-based grammar filtering approach due to the need to synthesize additional expressions in both ps and inv.
@@ 6.5 Comparison with LLMs
To test the capabilities of LLMs in generating code for new or low-resource DSLs, we prompt a state-of-the-art proprietary state of the art LLM Claude Opus [1] (our evaluation using other LLMs such as GPT4 shows similar results).
We selected three backends for this experiment: MLX, a completely new DSL, Gaudi, and Gemmini, which are not well represented in the training corpus of these models.
We prompted Claude Opus to generate code for these DSLs for the linear_dodge benchmark from the blend suite. The prompt instructions are shown in  Fig. 12. In Fig. 13 we show the code generated by the LLM for the three prompts.
Upon analysis, we found that all three generated programs were incorrect. The Gemmini-generated code in  Figure 13(a) partially uses the correct APIs (mvin, mvouts), but the computation with config_ex is incorrect. For the Gaudi-generated code in Figure 13(b), the model hallucinates the TPC library, which does not exist in the actual Gaudi programming model. The MLX-generated code in Figure 13(c) has the correct call to the library function add, but the imports are incorrect, making the code non-functional. In addition to the generated code being incorrect, it is challenging to verify these outputs formally as syntactically translating the generated code to SMT-LIB is not trivial. The experiment highlights two significant challenges in generating verified code using LLMs mentioned earlier. In contrast, Tenspiler, which uses a verified lifting-based approach, can easily handle these challenges.
@@
Figure 13: LLM generated code for the prompt
@@ 7 Related Work
Verified Lifting. Verified lifting uses program synthesis to translate code instead of designing traditional pattern-matching compilers, and has been used across application domains [15, 5, 6, 20]. Adapting these prior compilers for translation to tensor operations is nontrivial. Tenspiler introduces a novel tensor algebra-based to make synthesis efficient, and supports a diverse set of backends.
@@ Appendix A Backend Details
  • TPC-C for Intel Gaudi processor: Intel Gaudi processors are specialized accelerators for high-performance AI, featuring a Matrix Multiplication Engine for deep-learning operations and a programmable Tensor Processor Core (TPC) cluster, a VLIW SIMD processor, that accelerates remaining operations.
    The way to execute programs on TPC cluster is by writing those programs using TPC programming language. Intel’s SynapseAI TPC SDK includes an LLVM-based TPC-C compiler to compile those programs. Specifically, the TPC programming language, TPC-C in short, is a derivative of C99 with added data types that enable easy utilization of TPC’s SIMD capabilities. It natively supports wide vector data types to assist with programming of TPCs (for example, float64, uchar256, etc.)
    A TPC program consists of two parts: TPC kernel code and host glue code. TPC kernel code is the ISA executed by the TPC processor. The host code is executed on the host machine and specifies details of how the program input/outputs can be dynamically partitioned between the numerous TPC processors in the Gaudi processor.
@@
Table  1 shows the number of DSL APIs used in the kernel code generated by Tenspiler for all the target DSLs and for all the benchmarks used in the experiments. Note that we only count the number of DSL APIs of kernel code, not the host code (or the glue code) that is needed to launch the kernels on GPUs or accelerator devices.
Benchmark C++ Numpy TensorFlow PyTorch MLX TPC-C Gemmini
color_burn 18 5 5 5 5 21 NA
color_dodge 18 4 4 4 4 20 NA
darken_blend 18 2 2 2 2 4 NA
dissolve_blend 19 8 8 8 8 7 NA
lighten_blend 18 2 2 2 2 4 NA
linear_burn 14 2 2 2 2 6 2
linear_dodge 14 1 1 1 1 4 1
multiply_blend 14 2 2 2 2 4 NA
normal_blend_8 7 4 4 4 4 17 5
normal_blend_f 8 4 4 4 4 8 3
overlay_blend 18 12 12 12 12 32 NA
screen_blend 14 4 4 4 4 12 4
matmul 13 1 1 1 1 NA 1
rmsnorm1 6 2 2 2 2 NA 3
rmsnorm2 8 7 8 8 7 NA NA
softmax1 7 2 2 2 2 NA NA
softmax2 8 3 3 3 3 NA NA
softmax3 7 2 2 2 2 NA 2
softmax4 7 2 2 2 2 NA NA
transformer1 18 13 14 14 14 NA NA
transformer2 17 10 10 10 10 NA NA
transformer3 8 7 7 7 7 NA NA
transformer4 7 3 3 3 3 NA NA
Table 1: Lines of code for C++ inputs and the number of DSL APIs in code generated by Tenspiler for blend and Llama. (For the Llama benchmarks, we do not generate TPC-C code as most use matrix multiplication, which is not supported via TPC-C. For these benchmarks, we use PyTorch DSL for Gaudi processor.)
Appendix B Benchmark Name Mapping

### v3

@@ 5 Synthesis Optimizations
Figure 6: A depth 4 general synthesis grammar for the source in Figure 1(a).
@@ 5.2 Restricting Program States
We further optimize the search space by restricting the set of program states in Eq. 2. Instead of satisfying the VC for all $\sigma$, we find $PS$ and $inv$s that satisfy a bounded set. Bounded synthesis is crucial because most SyGuS solvers have limited support for recursive function definitions and require SMT solvers for validation. However, SMT solvers lack inherent support for reasoning about Tensir operators that are not covered by the standard theories defined in SMT-LIB [9] and require additional axioms to be defined. We instead integrate bounded synthesis by restricting the maximum unrollings of recursive operators, thereby eliminating the need for additional axioms. Specifically, we restrict the program states by limiting the lengths of the data structures and the sizes of the data types. For instance, in Tenspiler, we constrain all 1D tensors to length 2 and the integers to 6 bits or less for the first rounds of synthesis. In Fig. 6, all the tensor literals in $t_{1}$ are changed from an unbounded length “n” (shown in orange) to length 2 (shown in blue). If the synthesized choices fail to verify, we then increase the bounds in subsequent rounds. Note that since the synthesized solutions only work for a restricted set of program states, we invoke the theorem prover for subsequent verification to check if $PS$ and $inv$s are valid for all states.
@@ 5.3 Leveraging Expression Trees
Despite the above two optimizations, the synthesis search space remains large. For example, in the context of S in Figure 1(a) for which we need to synthesize two $inv$s and one $PS$, a depth-4 grammar, after removing the reduction operations, still presents around 100k potential solutions just for $PS$. Tensir plays a significant role in the further pruning of this search space. The design of Tensir operators effectively bridges the gap between high-level tensor operations and the loop-based paradigm commonly used for computing on tensors. This property of Tensir allows us to leverage the expression-tree-based filtering approach, which we describe next, to efficiently prune the synthesis search space.
@@ 5.4 Constraining Variables
The final optimization is to pinpoint specific variables (vars such as a,b in Figure 1(a)) and literals (lits such as 255 in Figure 1(a)) to be used in the grammar. Specifically, we constrain the variables to the set of live variables and also constrain constants to the set of constants that have appeared in the program. This strategy simplifies the computational task, avoiding the complexity of synthesizing a complete depth-4 operator sequence. By leveraging our expression tree-based approach, the search space reduces to 64 expressions, and the synthesizer promptly yields the correct solution within 76 secs: t_t(t_t(b, a, +), t_s(t_t(b, a, *), 255, /), -).
@@ 5.5 Overall Synthesis Algorithm
The algorithm described in Fig. 7 summarizes the synthesis phase in Tenspiler. This phase is used to search for the bodies of $PS$ and invariants which satisfy the VC. The synthesis is an iterative process conducted over multiple rounds, assuming a filtered search space leveraging type-based and expression tree optimizations described earlier. We start with the tensor bound size set to 2 which corresponds to restricting the program states optimization. In each round, we invoke Rosette’s search algorithm (line 5) to generate candidates for $PS$ and $inv$s. Upon obtaining a solution, the candidate undergoes validation against the VC for all program states, as the synthesis phase only checks within a constrained set of program states. We invoke a verifier (CVC5) (line 8) to perform this check. If the verifier yields “UNSAT,” the generated candidates are correct. Conversely, if “SAT” is returned, indicating incorrect candidates (line 11), the VC is augmented with blocking constraints. These constraints state that the generated $PS$ or $inv$s in next round should differ from those in the previous rounds. This iterative process continues for a specified number of rounds (max_rounds) before incrementing the tensor bound sizes. In cases where Rosette’s search algorithm does not produce a solution initially (line 13), indicating an overly restrictive grammar, the initial grammar is expanded to include additional options for both $PS$ and invariants, such as choices for loop bounds, indexing, and operator sequences. We keep a separate timer (not shown in Fig. 7) that maintains a maximum time bound for the entire synthesis process.
1def synthesis_algorithm(spec, tensor_size_bound, holing_grammar, search_algorithm, verifier, max_rounds, timeout):
2 r = 0 # rounds within one list bound
@@
8 if verifier(specification, ps_r, inv_r) == "UNSAT": return ps_r, inv_r
@@
14 # Increment tensor size bound
15 return synthesis_algorithm(spec, tensor_size_bound + 1, holing_grammar, search_algorithm, verifier, max_rounds, timeout)
@@ 6 Experiments
    simpl_array: 5 functions for computations on integer arrays originally from prior work [35].
@@ 6.1.1 Datasets for Evaluation
For the Llama benchmarks, we evaluate the synthesized code using weights from Vicuna [16], an open-source LLM with similar model size as Llama2.666We did not use Llama2 model weights as they are not publicly available. Some kernels operate over model weights, for which we directly use the weight matrices from Vicuna. For kernels operating over inputs, we simulate embeddings by creating random 32-bit float vectors within the range [0, 1). The evaluation primarily uses the 33B-parameter Vicuna model; however, for evaluating the MLX framework, the 7B-parameter version is used due to memory limitations.
@@ 6.1.2 Target Software Frameworks and Hardware Accelerators
[…]
@@ 6.2 Synthesis Timings
In this section, we present the time Tenspiler takes to synthesize and verify solutions for each of our benchmarks. During the synthesis phase, we apply all the optimizations mentioned in Sec. 5 with a 1 hour timeout. Tenspiler synthesizes the correct translations for all the benchmarks under 15 mins.
(a) blend benchmarks (image processing).
(b) utdsp benchmarks.
(c) DSP benchmarks.
(d) simpl_array benchmarks.
(e) darknet benchmarks.
(f) blas benchmarks.
(g) mathfu benchmarks.
(h) dspstone benchmarks.
(i) Llama benchmarks (ML kernels).
(j) makespeare benchmarks.
Figure 8: Synthesis timings for all the benchmark suites. Benchmark name legend in Sec. 10 in the Appendix.
@@ 6.2.1 Analysis
We observe that synthesis difficulty is correlated to both the number of loops and the complexity of the Tensir solution. For example, the dot benchmark in the blas benchmark set has a single loop. Its Tensir solution is reduce_sum(t_t(a, b, *)), which has two operators and only two arguments, a and b, to be synthesized. This benchmark synthesizes in around two seconds. On the other hand, the transformer_part1 benchmark from the LLama benchmark set has a doubly-nested loop and a complex solution with six operators. All arguments to these operators must be synthesized, with some requiring 3 expressions such as head * head_size + head_size and sqrt(head_size * 1). This benchmark takes around 1300 seconds to synthesize.
In addition to synthesis challenges, we recognize that the tree approach may restrict the ability of Tenspiler to synthesize different solutions. However, through manual verification, we confirm that Tenspiler consistently generates optimal solutions across our benchmarks. We evaluate the solutions using expression length as the cost function. Generally, shorter expressions mean fewer function calls and thus lower execution cost. To illustrate, we use the linear_dodge example from the blend benchmarks for which the synthesized solution is as follows:
@@
Figure 9: Kernel speedup over baseline. Benchmark name legend in Sec. 10 in the Appendix.
@@ 6.3 Performance Timings
Figure 10: E2E speedup over baseline. Benchmark name legend in Sec. 10 in the Appendix.
@@
1vector<float> matmul(vector<vector<float>> weight, vector<float> input) {
2 vector<float> output;
3 int m = weight.size();
4 int n = input.size();
5 for (int row = 0; row < m; row++) {
6 float curr = 0;
7 for (int col = 0; col < n; col++) {
8 curr += weight[row][col] * input[col];}
9 output.push_back(curr);}
10 return output;}
(a) Original matmul function in C++.
1@cuda.jit()
2def matmul (weight, input, res):
3 m = len(weight)
4 n = len(input)
5 for i in range(m):
6 curr = 0
7 for j in range(n):
8 curr += weight[i][j] * input[j]
9 res[i] = curr
(b) Numba kernel annotated version of matmul.
Figure 11: Manually rewritten Numba example.
Compare Against Pattern Matching-Based Compilers. As outlined in Sec. 1, Tenspiler is designed to address the limitations inherent in traditional transpilers that rely on pattern matching to compile. Such compilers are resource-intensive to develop and prone to errors. To the best of our knowledge, no existing compiler matches the breadth of DSL support offered by Tenspiler. However, specialized compilers, such as Numba [29], have been introduced for accelerators like GPUs. Numba leverages LLVM IR to generate GPU-accelerated code from Python code, making it a suitable candidate for comparison.
For benchmarking purposes, we utilize the same datasets, test cases, and setup described previously in Sec. 6. Benchmarks are rewritten in Python and adapted to conform to CUDA kernel requirements by removing return statements, as shown in Fig. 11. Additionally, relevant data are cast to NumPy arrays as Numba focuses on optimizing code written against NumPy’s API. These syntactic requirements represent a limitation of Numba’s approach. In contrast, Tenspiler operates directly on the original benchmark implementations.
Experimental results demonstrate that GPU-based PyTorch and TensorFlow code generated by Tenspiler performs, on average, 1.87$\times$ faster than code annotated with Numba. Remarkably, while Numba benefits from years of development by expert engineers, Tenspiler achieves superior performance with only 200 additional lines of code dedicated to code generation. A closer examination of the compiled PTX assembly code for the matmul benchmark, which shows a 2.6$\times$ speedup, reveals that the Numba-generated code lacks the use of advanced instructions and techniques such as fused multiply-add (FMA), tiled-based computation models, or shared memory, [^9] See Sec. 12 in the Appendix for the PTX code. which are crucial for peak performance. These techniques are standards in PyTorch and TensorFlow with optimized kernels. In contrast, Numba requires extensive manual tuning to implement, evident in the more complex and faster matmul example in its documentation. [^10] For the detailed example of an optimized matmul function with shared memory for Numba, see https://numba.readthedocs.io/en/stable/cuda/examples.html#id30 Tenspiler, by automatically recognizing and translating matrix multiplication operations to leverage the pre-optimized kernels, avoids the complexities of manual code optimization while achieving high performance.
@@ 6.4 Ablation Study
Figure 12: Synthesis timings for all the benchmarks with and without Tenspiler’s tree-based optimization. Benchmark name legend in Sec. 10 in the Appendix.
@@
As illustrated in Fig. 12, without the tree based optimization, 42 out of the total 69 benchmarks timed out. In particular, all benchmarks of the blend, blas, dspstone, makespeare, and utdsp suites timed out. For the benchmarks that succeed, the synthesis phase slows down by an average 101.55$\times$ compared to the tree-based grammar filtering approach due to the need to synthesize additional expressions in both $PS$ and $inv$s.
@@ 6.5 Comparison with LLMs
To test the capabilities of LLMs in generating code for new or low-resource DSLs, we prompt a state-of-the-art proprietary LLM Claude Opus [1] (our evaluation using other LLMs such as GPT4 shows similar results). We selected three backends for this experiment: MLX, a completely new DSL, Gaudi, and Gemmini, which are not well represented in the training corpus of these models. We prompted Claude Opus to generate code for these DSLs for the linear_dodge benchmark from the blend suite. The prompt instructions are shown in  Fig. 13. In Fig. 14 we show the code generated by the LLM for the three prompts.
[…]
@@
Figure 14: LLM generated code for the prompt.
6.6 Extension to Higher-Dimensional Tensors
Our benchmarks only involve 1D and 2D tensors, as most operations are performed on images (the blend benchmarks) and weight matrices (the LLama benchmarks). In this section, we demonstrate that Tenspiler can be extended to support higher-dimensional tensors with the generalizability of Tensir and the synthesis optimizations discussed in Sec. 5. Specifically, we extend Tensir to accommodate 3D tensors and all corresponding element-wise operations. Additionally, we adapt the operator restriction optimization (introduced in Section 5.1) to apply to 3D tensors. When the source program returns a 3D tensor, our grammar is restricted to include only element-wise 3D tensor operations. We also retain the program state restriction optimization technique from Section 5.2. Furthermore, we extend our support to leverage expression trees performed on individual elements in tensors, as detailed in Section 5.3, to guide the search for vectorized operations within 3D tensor spaces.
We evaluate Tenspiler’s synthesis optimizations on artificial benchmarks involving 3D tensors. We create these benchmarks by combining random element-wise operations. The maximum depth of these benchmarks is chosen to be 5 to match that of all our existing real-world benchmarks, as described in Sec. 6. Results in Fig. 15 show that the synthesis time grows linearly with the depths of the benchmarks. The depth-1 benchmark synthesizes the fastest in 6 seconds, while the depth-5 benchmark takes the longest, in 184 seconds.
The sharp increase in timing for depth 5 expressions in Fig. 15 is due to the number of expressions we are synthesizing and their complexity. A benchmark with 3 loops involves synthesizing 3 invariants and 1 post-condition, each with expression sizes up to depth 5. Despite these challenges, we easily extend Tenspiler’s optimizations and synthesize these benchmarks well within the 1 hour timeout. As future work, to further scale Tenspiler’s synthesis algorithm for handling more complex benchmarks, we could explore strategies such as guiding the search process using machine learning techniques, implementing bottom-up synthesis starting with inner loops first, performing bounded synthesis with unrolled loops, and combining these approaches with Tenspiler’s current synthesis optimizations.
Figure 15: Synthesis timings for artificial 3D tensor benchmarks.
@@ 7 Related Work
Verified Lifting. Verified lifting uses program synthesis to translate code instead of designing traditional pattern-matching compilers, and has been used across application domains [15, 5, 6, 20, 22]. Adapting these prior compilers for translation to tensor operations is nontrivial. Tenspiler introduces a novel tensor algebra-based to make synthesis efficient, and supports a diverse set of backends.
@@ 9 Backend Details
[…]
    for Intel Gaudi processor: Intel Gaudi processors are specialized accelerators for high-performance AI, featuring a Matrix Multiplication Engine for deep-learning operations and a programmable Tensor Processor Core (TPC) cluster, a VLIW SIMD processor, that accelerates remaining operations.
[…]
@@
10 Benchmark Name Mapping